A stone is thrown from a sling at a speed of 40m/s at an angle of 50 °. Find (a) time of ascent; (b) maximum height reached; (c) range.
1
Expert's answer
2021-10-15T10:23:13-0400

(a) The time of ascent is


"t=v\sin\theta/g=3.13\text{ s}."


(b) Maximum height reached:


"H=(v\sin\theta)^2/2g=47.9\text{ m}."


(c) Range:


"R=v^2\sin(2\theta)/g=161\text{ m}."